OpenStudy (allieatthedisco):

Looking for help on how to find the slope-intercept form that is perpendicular to the given line. 4x-2y=3 (2,1)

OpenStudy (helder_edwin):

u have \[\large 4x-2y=3 \] which can be rewritten as \[\large 2y=4x-3\qquad\text{or}\qquad y=2x-\frac{3}{2} \] this is the slope-intercept form. this line has slope m=2

OpenStudy (helder_edwin):

now, a line perpendicular to the given one has slope m1 that satisfies: \[\large -1=m\cdot m_1=2m_1\qquad\text{so}\qquad m_1=-\frac{1}{2} \] so it slope-intercept equation would be: \[\large y=-\frac{1}{2}x+b\tag{*}\]

OpenStudy (helder_edwin):

finaly, since this line has to pass through (2,1) replace x=2 and y=1 in (*) to get the value of b: \[\large 1=-\frac{1}{2}\cdot2+b\qquad\Rightarrow\qquad b=2 \] So the equation u r looking for is \[ \large y=-\frac{1}{2}x+2 \]